OpenStudy (anonymous):

The formula for determining interest compounded monthly is A = P(1 +r/12 )^12t, where A represents the amount invested after t years, P the principal invested, and r the interest rate. Jimmy invests $1,000 at an interest rate of 10% for 3 years, while Jenny invests $1,000 at an interest rate of 5% for 6 years. Determine the amount of return gained by Jimmy and Jenny. In complete sentences, summarize your results.

OpenStudy (anonymous):

You need to use the formula to work out the amount for Jenny and Jimmy first. Do you have the answers to this?

OpenStudy (anonymous):

Also interest gained is I=A-P where A is the final amount and P is the principle or amount first invested.

OpenStudy (anonymous):

I do not have those answers.

OpenStudy (anonymous):

Jimmy: A = $1,000 ( 1 + .1/12 ) ^36 = $1,346.18 Jenny: A = $1,000 ( 1 + .05/12 ) ^72 = $4,413.15

OpenStudy (anonymous):

Jimmy gains: $1,346.18 - 1,000 = $346.18 Jenny gains: $4,413.15 - 1,000 = $3,413.15

OpenStudy (anonymous):

Well since Chlorophyll has given you the answers what can you say about the results?

OpenStudy (anonymous):

@claireshowers say something!

OpenStudy (anonymous):

@claireshowers Don't forget to write the conclusion, It's for your own benefit!